OpenStudy (anonymous):

simplify by removing factors of 1 n^2-16/(n+4)^4

jimthompson5910 (jim_thompson5910):

Hint: n^2-16 factors to (n-4)(n+4) using the difference of squares rule

jimthompson5910 (jim_thompson5910):

so what does that mean?

OpenStudy (anonymous):

huh i dont understand

jimthompson5910 (jim_thompson5910):

If I have n+4 up top and a n+4 down on the bottom, what will happen?

OpenStudy (anonymous):

that it would be n-4 right

OpenStudy (anonymous):

it would then be n+1

jimthompson5910 (jim_thompson5910):

yes, the n+4 up top cancels with one of the n+4 terms down on the bottom

OpenStudy (anonymous):

or n+8

jimthompson5910 (jim_thompson5910):

not sure where you're getting n+1

OpenStudy (anonymous):

so my answer then would be n+4

jimthompson5910 (jim_thompson5910):

not quite, let me draw it out for you

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

alright

jimthompson5910 (jim_thompson5910):

\[\Large \frac{n^2-16}{(n+4)^4}\] Note: in step 2, i'm factoring using the difference of squares law \[\Large \frac{(n-4)(n+4)}{(n+4)^4}\] \[\Large \frac{(n-4)(n+4)}{(n+4)(n+4)(n+4)(n+4)}\] \[\Large \frac{(n-4)\cancel{(n+4)}}{\cancel{(n+4)}(n+4)(n+4)(n+4)}\] \[\Large \frac{n-4}{(n+4)(n+4)(n+4)}\] \[\Large \frac{n-4}{(n+4)^3}\] So \[\Large \frac{n^2-16}{(n+4)^4}\] completely simplifies to \[\Large \frac{n-4}{(n+4)^3}\] In other words, \[\Large \frac{n^2-16}{(n+4)^4} = \frac{n-4}{(n+4)^3}\]

jimthompson5910 (jim_thompson5910):

let me know if you have questions

OpenStudy (anonymous):

so then i would also subtract as well and so my answer then would be n-4/(n+4)^3 right

jimthompson5910 (jim_thompson5910):

yes n-4 all over (n+4)^3